OpenStudy (anonymous):

Multiply (4 + i)(1 – 5i)

OpenStudy (anonymous):

4 - 4i - 5i^2

OpenStudy (anonymous):

\[(4 + i)(1 – 5i) = 4 - 20i + i + 5 = 9 - 19i\]

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

may you help me on another problem

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

\[rationalize the denominator 9/(7-3\sqrt{3})-(6-4\sqrt{3})\]

OpenStudy (anonymous):

Is all of (7−3√3)−(6−4√3) under the 9?